Diana has been saving nickels and dimes. She opened up her piggy bank and determined that it contained 48 coins worth $4.50. Determine how many nickels and dimes were in the piggy bank.
Answer by mananth(16949)

Diana has been saving nickels and dimes. She opened up her piggy bank and determined that it contained 48 coins worth $4.50. Determine how many nickels and dimes were in the piggy bank.
...
nickels -----x
dimes -------y
x+y = 48........................1
..
5x+10y=450......................2
multiply (1) by -5
-5x-5y=-240
add this to (2)
5y=210
/5
y=42 dimes
..
48-42= 6 nickels
...
m.ananth@hotmail.ca